You are considering buying a new car. The sticker price is $37,750 and you have $5,000 to put toward a down payment. If you can negotiate a Flat interest rate of 7.5 percent and you wish to pay for the car over a 4-year period. What is the equivalent Annual Percentage Rate on this loan?
Answer #1

Balance payment = 37750-5000= 32750

Total interest payment= 32750*7.5%

= 2456.25.

Annual interest payment= 2456.25/4= 614.0625

Equivalent annual percentage = 614.0625/32750

= 1.875%